OpenStudy (anonymous):

Solving Quadratic By Factoring 1.) X^2+5x=-6 2.) 7r^2-14r=-7 3.) 5r^2-44r+120=-30+11r

OpenStudy (ranga):

For each of the problem make the right hand side zero by subtracting the right hand side terms from both the left and the right. Simplify and then factor them. 1) x^2 + 5x + 6 = 0 x^2 + 3x + 2x + 6 = 0 x(x + 3) + 2(x + 3) (x + 2)(x + 3) = 0 x = -2 or x = -3 For the second problem you factor out 7 and it will be easier to factor.